# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ VII Generalization to two dimensions: group loop models
where $|r_{i}|\leq 3$ for all $i$; this holds from the fact that any group (not semigroup) exhibits a finite presentation satisfying $|r_{i}|\leq 3$ (see App. A for the proof). Writing $r_{i}=\mathtt{g}_{m}\mathtt{g}_{n}\mathtt{g}_{\ell}$, this corresponds to an object which we refer to as a net:
@@
and so on. Any loop configuration corresponding to one of the above nets can be created or destroyed without changing $\varphi[w(\gamma)]$ (the group element associated to $w(\gamma)$), since for any curve $\gamma$ that cuts across the net, creating or destroying the net simply corresponds to applying the appropriate relation $r_{i}$ at some point in the word $w(\gamma)$. The dynamics we consider are generic dynamical processes which preserve $\varphi[w(\gamma)]$ for all closed curves $\gamma$, which may be viewed as an unusual type of gauge constraint. Thus the dynamics will include processes which nucleate and destroy nets associated to each $r_{i}$, and will also include processes where lines are moved, stretched, contracted, and where intersections of lines are moved. It will also include processes which attach a loop with an intersection point of other loops, like so:
@@
The dynamical processes above were formulated for the case where $G$ is a group, but it is also possible to generalize to the case when $G$ is merely a semigroup, producing models reminiscent of the 2D generalizations of the Motzkin chain studied in Ref. Balasubramanian et al. (2023). Obtaining a more systematic understanding of the temporal and spatial complexity of the dynamics of these models would constitute an interesting avenue for future work. One final observation we make is that there is a simple way to map group loop models to what we call tile models. Observe that a configuration of loops or nets splits the plane into a set of disjoint tiles (two adjacent tiles are joined by an edge which corresponds to a generator of the group). Then, it turns out to be possible to label tiles with an element $g\in G$ such that if two tiles have labels $g$ and $h$ and share an edge corresponding to generator $\mathtt{k}$, then $g\mathtt{k}=h$. For example, if $G=\mathbb{Z}$, each tile will be labelled by an integer and two adjacent tiles have labels which differ by $1$ (which is the generator of $\mathbb{Z}$). This corresponds to a mapping to a height model, of which the tile model is a generalization of in the case of arbitrary $G$. This mapping may play a role in understanding the nature of fragmentation in these models, which we leave to future work.
@@ VIII Discussion
In this work, we have constructed a number of natural dynamical systems—with local few-body interactions—in which relaxation places anomalously expensive demands on a system’s temporal and/or spatial resources. When the models have local conserved densities, the hydrodynamics of these densities is anomalous or frozen; even when conserved densities are absent, we have presented diagnostics for nonergodic behavior. Our examples were all constructed in the context of models with intrinsic Hilbert space fragmentation. A natural question is whether the intrinsic fragmentation is essential to their physics. In our framework, dynamics without fragmentation is generated by finite presentations of the trivial group, which cannot have a superlinear Dehn function (see App. B). This does not mean that the dynamics of these models cannot be slow, but it does mean that any mechanism for slow thermalization must originate from something other than the Dehn function. Our results lend themselves to several natural extensions. Most naturally, the anomalous hydrodynamic relaxation we saw in the ${\sf BS}$ model can be extended to other groups with presentations that manifest a conservation law. Whether these groups give rise to new classes of hydrodynamic relaxation is an interesting question for future work (one such example will be presented in Balasubramanian et al. (2024)). Another interesting direction is to investigate the ground states of Hamiltonians that implement group dynamics. A natural class of frustration-free Hamiltonians can be read off from the transfer matrices of bistochastic Markov processes Rokhsar and Kivelson (1988); their ground states are the equal-weight superpositions of all the configurations in a sector, and their spectral gaps can be bounded by the Markov-chain gap. The tools developed here are also useful for understanding these ground-state and spectral properties. The family of models we considered is restricted in the sense that the dynamical constraints can be expressed in the computational basis, in that every computational-basis product state is in a definite dynamical sector of Hilbert space. More generally, one can consider constraints associated with a commuting set of projectors with entangled eigenstates. In one-dimensional spin chains, such commuting projectors can be deformed into unentangled projectors by a short-depth unitary circuit. However, this process just corresponds to conjugating the Hamiltonians/unitaries we have explored with short-depth unitary circuits, and yields nothing qualitatively new. Getting something new in one dimension thus requires defining constraints using non-commuting projectors—as occurs for example in the (quantum-fragmented) Temperley-Lieb model Batchelor and Kuniba (1991); Aufgebauer and Klümper (2010); Moudgalya and Motrunich (2022)—and hence pursuing this direction necessitates developing a more systematic understanding of such constraints. In higher dimensions however, there are sets of commuting projectors (like those of the toric code) whose simultaneous eigenstates are inherently long-range entangled. An interesting task for future work would be to extend our group-theoretic dynamical systems to these models, and explore the resulting dynamics of entanglement. Finally, it would be interesting to explore the stability of our results with respect to weak violations of the constraints. One way to do this is to break the constraints in an isolated region of space, or by bringing a thermal bath in contact with the system at its boundary. In the case of group dynamics, doing so generically destroys fragmentation, making the dynamics ergodic. It furthermore leads to all states being connected by at most $O(L^{2})$ steps of the resulting dynamics, since the $O(L)$ elements of a given product state’s geodesic word need to be moved at most an $O(L)$ distance to the constraint-free region, at which point they can be changed into the elements of any other geodesic word. It is however still possible for the dynamics in this case to have long thermalization times, due to bottlenecks in Fock space that arise from the fact that the constraints are only broken in a single region of space. A specific example of this will be presented in Han et al. (2023).
@@ IX Acknowledgements
Appendix A Group presentations
In this section we review some basic notions about discrete group presentations and prove a few small results mentioned in the main text. Formally, a discrete group is a set equipped with a binary associative operation under which one element acts an the identity, and is such that each element of the set has a corresponding inverse. It is common to discuss a group $G$ in terms of a specific set of generators $S$ and relations $R$ between them, writing
$G=\langle S\,|\,R\rangle$ (114)
to signify this relationship. For example, one might usually think of the group $\mathbb{Z}^{2}$ as being defined by two generators $S=\{\mathtt{x},\mathtt{y}\}$ that obey the single non-trivial relation $\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}$ (this being in addition to trivial relations like $\mathtt{x}\mathtt{x}^{-1}=\mathtt{e}$). [^24] We will always use multiplicative notation for the group operation, even for Abelian groups. This in some respects is however too narrow of a viewpoint, since it is possible for different choices of $S$ and $R$ to produce the same group. As an example, consider the groups
@@ Appendix A Group presentations
$G_{2}$ $=\langle\mathtt{x},\mathtt{y}\,|\,\mathtt{x}^{3}=\mathtt{y}^{2}\rangle$
It is easy to see that $G_{1}\cong G_{2}\cong\mathbb{Z}$, despite $G_{1}$ and $G_{2}$ having a different number of generators and relations. A given group in general admits an infinite number of different presentations, but below we will prove that the group theoretic functions defined in the main text—the Dehn function, expansion length, and so on—have asymptotic scaling behaviors that are independent of the choice of presentation. We can exploit this fact to choose presentations satisfying some particular desired property. For example, in the context of group dynamics, we may be concerned with choosing a model of dynamics where the Hamiltonian or unitary gates are as local as possible. Since the locality of the dynamics is limited by the maximum size of the relations in $R$, we would thus like to minimize the size of the relations. To this end we have the following proposition:
@@
Every finitely generated group has a presentation $\langle S|R\rangle$ where all relations $r_{i}\in R$ are have length $|r_{i}|\leq 3$.
@@
Consider the Cayley 2-complex ${\mathsf{CG}}_{G}$ of a finitely presentable group $G$ (see Sec. IV for a brief discussion of its definition) obtained from a finite presentation $P$. While the exact structure of ${\mathsf{CG}}_{G}$ depends on $P$, ${\mathsf{CG}}_{G}$ can always be subdivided to obtain a simplicial complex where each 2-cell has 3 edges. Since each 2-cell in the complex corresponds to a relation, $|r_{i}|\leq 3$ for all $r_{i}\in R$ in the subdivided complex, thereby defining a presentation $P^{\prime}$ whose relations all have length $\leq 3$. Since $P$ was assumed finite this subdivision is completed after only a finite number of steps, and $P^{\prime}$ is consequently also finite. ∎
Note that while the Cayley 2-complex of a semigroup can also be subdivided, the lack of translation invariance in a semigroup’s Cayley complex means that the resulting subdivision may yield a presentation with infinitely many generators (an illustrative example in this regard is to compare the semigroup $\mathtt{n}\times\mathtt{n}$ with the group $\mathbb{Z}\times\mathbb{Z}$). As mentioned above, the group theoretic properties we are interested in from the point of view of group dynamics are largely insensitive to the choice of presentation. For example, recall the growth function $N_{K}(L)\triangleq|\{g,\,|g|\leq L\}|$ defined in (28) of the main text, which measures the number of dynamical sectors as a function of system size. The scaling of $N_{K}(L)$ with $L$ is independent of the choice of presentation, allowing us to meaningfully talk about the growth function of a group, rather than of a presentation:
@@
Let $N_{K,P}$ denote the growth function for a particular presentation $P=\langle S|R\rangle$ of $G$. Then
@@
Let $P=\langle a_{1},\dots,a_{|S|}\,|\,R\rangle$ and $P^{\prime}=\langle a_{1}^{\prime},\dots,a^{\prime}_{|S^{\prime}|}\,|\,R^{\prime}\rangle$. Then since both $P,P^{\prime}$ present $G$, each $a_{i}^{\prime}$ can be expressed as a product of a finite number of $a_{i}$. Let $n_{PP^{\prime}}$ denote the maximal number of generators of $P$ that appear when writing the $a_{i}^{\prime}$ in terms of these generators. Let also $|g|_{P}$ denote the geodesic distance of $g\in G$ with respect to the presentation $P$. Then $|g|_{P^{\prime}}\leq n_{PP^{\prime}}|g|_{P}$. Thus
@@
By symmetry, we may also perform a similar rewriting of the generators of $P$. Thus there exist $O(1)$ constants $n_{PP^{\prime}},n_{P^{\prime}P}$ such that
@@
Similar reasoning can be applied to show that the Dehn function and expansion length (see Sec. II or the following appendix for definitions) of a group are presentation-independent:
@@
Let ${\sf Dehn}_{P}(L)$, ${\sf EL}_{P}(L)$ be the Dehn function and expansion length of a group with a finite presentation $P=\langle S|R\rangle$. Then if $P,P^{\prime}$ are any two such finite presentations,
@@ Appendix B A primer in geometric group theory
In this appendix we state and prove some useful facts about the geometry and complexity of finitely presentable discrete groups. Many of the statements derived below are well-known results in the math literature, and we have tried to provide citations to the original works when appropriate. A particularly accessible review of background material relevant to the discussion to follow can be found in Ref. Clay and Margalit (2017); a more advanced reference is Ref. Sapir (2011). We will mostly be interested in infinite groups, since finite ones have trivial large-scale geometry (in a sense soon to be made precise). Finitely presentable infinite groups are of course very easily constructed; indeed it is easily verified any presentation where the number of generators exceeds the number of nontrivial relations [^25] By a “trivial relation” we mean a free reduction / expansion, viz. a relation of the form $\mathtt{a}\mathtt{a}^{-1}=\mathtt{e}$. will generate an infinite group. Free groups (on $n>1$ generators) and Abelian groups in some sense define opposite geometric extremes, and most of the interesting action will happen in cases where an intermediate amount of Abelian-ness is introduced to a non-Abelian free group.
@@
The definition of the Dehn function (14) in the main text relates only to the (worst-case) complexity of deforming a given word $w\in K_{e}$ into the identity word (recall that $K_{g,L}$ is the set of length-$L$ words that represent the element $g$, viz. $K_{g}=\{w\,|\,|w|=L,g(w)=g\}$). We claimed in the main text that focusing on the complexity of words in $K_{e}$—as opposed to $K_{g}$ for $g\neq e$—was done without loss of generality. We now prove that looking at $K_{g\neq e}$ does indeed not produce anything which is not already captured by ${\sf Dehn}(L)$:
@@
up to unimportant contributions linear in $L$ [^26] They are “unimportant” because we need $O(L)$ steps just to get rid of trivial words like $ww^{-1}$. However, an extra $L$ steps can be very important in certain settings, especially where e.g. we do not allow words of length greater than $L$ to appear at any stage of the deformation (which we very well may want to do on physical grounds!). This means that ${\sf Dehn}(L)\lesssim{\sf Dehn}_{g}(L)$. Conversely, since we can deform $e$ to $w_{1}^{-1}w_{2}\sim e$ in time $\sim{\sf Dehn}(w_{1}w_{2}^{-1})$, $w_{1}$ can be deformed into $w_{1}(w_{1}^{-1}w_{2})=w_{2}$ in time $\lesssim{\sf Dehn}(L)$. Thus we also have ${\sf Dehn}(w,w^{\prime})\lesssim{\sf Dehn}(w_{1}w_{2}^{-1})$, so up to factors of order $L$, we have
@@
The above argument means that—if the amount of space available to us is not restricted—all sectors have the same worst-case complexity. If we however require that all words involved have length $\leq L$, Prop. 4 is modified to read
@@
In this case, all sectors $K_{g,L}$ for which the geodesic length [^27] Recall that the geodesic length of a group element is defined by the length of the shortest word representing $g$, viz. $|g|=\min\{w\,:\,\varphi(w)=g\}$. of $g$ satisfies $|g|/L<1$ in the $L\rightarrow\infty$ limit will still have ${\sf Dehn}_{g}(L)\sim{\sf Dehn}(L)$. On the other hand, sectors where $|g|/L=1$ as $L\rightarrow\infty$ will not have enough “free space” for the above argument to work, and will thus have ${\sf Dehn}_{g}(L)<{\sf Dehn}(L)$ (in the case where fragile fragmentation occurs [see Sec. V], this remains true provided we define ${\sf Dehn}(w,w^{\prime})=0$ if $w_{1},w_{2}$ are in different subsectors of $K_{g,L}$). Sectors where $|g|/L=1$ as $L\rightarrow\infty$ are however necessarily exponentially smaller than those with $|g|/L<1$ (see Prop. 10), and hence a random state will be in a sector with ${\sf Dehn}_{g}(L)\sim{\sf Dehn}(L)$ with high probability. We now come to the topic of determining which sorts of groups we can expect to have large Dehn functions. A basic result is that only infinite non-Abelian groups can have interesting Dehn functions, due to the following easily verified statements:
@@
Free groups have ${\sf Dehn}(L)\sim L$.
@@
This follows from the fact that the Cayley graphs of free groups are trees, and thus any word $w\in K_{e}$ must represent a path which retraces its steps at every point and forms no loops. Such paths can always be contracted to the trivial path in $O(L)$ steps. [^30] For the discussion in this appendix, we are working in the convention where ${\sf Dehn}(w)$ is the number of relations needed to reduce $w$ to the identity $\mathtt{e}$, allowing the length of the word to change in the process (so that e.g. $\mathtt{e}\mathtt{e}$ may be replaced by $\mathtt{e}$). For physical models of group dynamics, where the total length of the word is always fixed, ${\sf Dehn}(w)$ will generically be longer by a factor of $L$, corresponding to moves of the form $\mathtt{e}\mathtt{a}=\mathtt{a}\mathtt{e}$ that need to be performed during the reduction of $w$ to $\mathtt{e}^{L}$.
To look for interesting Dehn functions we thus need infinite groups with a moderate amount of Abelianness, which (roughly speaking) possess nontrivial loops at all length scales. First examples of such groups are provided by those where ${\sf Dehn}(L)$ scales as a nontrivial polynomial. These turn out to be precisely those groups which are virtually nilpotent, as proven in Ref. Gersten et al. (2003). The simplest example is:
@@
has ${\sf Dehn}(L)\sim L^{3}$ Gersten et al. (2003).
The group ${\sf BS}(1,2)$ Baumslag and Solitar (1962) studied in detail in the main text is the simplest group with exponential Dehn function, whose properties we discuss in detail in App. E.
@@
As discussed in the main text, the space complexity of the word problem is given by the maximal size of words that one must encounter when reducing a $w\in K_{e}$ to the identity, as measured by the expansion length ${\sf EL}(w)$, a quantity originally introduced by Gromov in Ref. Gromov (1992). If ${\sf EL}(w)>|w|$, then $w$ must expand by a nontrivial amount while being reduced to the identity. As we did with the area, we may use the expansion length to define a distance between any two words $w_{1,2}$ that represent the same group element. Instead of doing something like ${\sf EL}(w_{1}w_{2}^{-1})$ (which we do not do on account of the homotopy relating $w_{1}$ to $w_{2}$ needing to be properly based), we define the relative expansion length ${\sf EL}(w,w^{\prime})$ between two words $w_{1}\sim w_{2}$ by replacing $e$ by $w_{2}$ in the above definition:

### 11 Apr 2024 v2

@@ VII Generalizations to two dimensions: group loop models
where each of the $r_{i}$ are words to be identified with the identity in $G$, and $|r_{i}|\leq 3$ for all $i$; this restriction on the length of the relations follows from the fact that any group (but not any semigroup) exhibits a finite presentation satisfying $|r_{i}|\leq 3$ (see App. A for the proof). Writing $r_{i}=\mathtt{g}_{m}\mathtt{g}_{n}\mathtt{g}_{\ell}$, this corresponds to an object which we refer to as a net:
@@
and so on. Any loop configuration corresponding to one of the above nets can be created or destroyed without changing $\varphi(w(\gamma))$ (the group element associated to $w(\gamma)$), since for any curve $\gamma$ that cuts across the net, creating or destroying the net simply corresponds to applying the appropriate relation $r_{i}$ at some point in the word $w(\gamma)$. The dynamics we consider are generic dynamical processes which preserve $\varphi(w(\gamma))$ for all closed curves $\gamma$, which may be viewed as an unusual type of gauge constraint. Thus the dynamics will include processes which nucleate and destroy nets associated to each $r_{i}$, and will also include processes where lines are moved, stretched, contracted, and where intersections of lines are moved. It will also include processes which attach a loop with an intersection point of other loops, like so:
@@
The dynamical processes above were formulated for the case where $G$ is a group, but it is also possible to generalize to the case when $G$ is merely a semigroup, producing models reminiscent of the 2D generalizations of the Motzkin chain studied in Ref. [61]. Obtaining a more systematic understanding of the temporal and spatial complexity of the dynamics of these models would constitute an interesting avenue for future work. One final observation we make is that there is a simple way to map group loop models to what we call tile models. Observe that a configuration of loops or nets splits the plane into a set of disjoint tiles (two adjacent tiles are joined by an edge which corresponds to a generator of the group). Then, it turns out to be possible to label tiles with an element $g\in G$ such that if two tiles have labels $g$ and $h$ and share an edge corresponding to generator $\mathtt{k}$, then $g\mathtt{k}=h$. For example, if $G=\mathbb{Z}$, each tile will be labelled by an integer and two adjacent tiles have labels which differ by $1$ (which is the generator of $\mathbb{Z}$). This corresponds to a mapping to a height model, of which the tile model is a generalization of in the case of arbitrary $G$. This mapping may play a role in understanding the nature of fragmentation in these models, which we leave to future work.
@@ VIII Discussion
In this work, we have constructed a number of natural dynamical systems—with local few-body interactions—in which relaxation places anomalously expensive demands on a system’s temporal and/or spatial resources. When the models have local conserved densities, the hydrodynamics of these densities is anomalous or frozen; even when conserved densities are absent, we have presented diagnostics for nonergodic behavior. Our examples were all constructed in the context of models with intrinsic Hilbert space fragmentation. A natural question is whether the intrinsic fragmentation is essential to their physics. In our framework, dynamics without fragmentation is generated by finite presentations of the trivial group, which cannot have a superlinear Dehn function (see App. B). This of course does not mean that the dynamics of models without fragmentation cannot be slow, but it does mean that any mechanism for slow thermalization must originate from something other than the Dehn function. Our results lend themselves to several natural extensions. Most naturally, the anomalous hydrodynamic relaxation we saw in the ${\sf BS}$ model can be extended to other groups with presentations that manifest a conservation law. Whether these groups give rise to new classes of hydrodynamic relaxation is an interesting question for future work (a family of such examples will be presented in [62]). Another interesting direction is to investigate the ground states of Hamiltonians that implement group dynamics. A natural class of frustration-free Hamiltonians can be read off from the transfer matrices of bistochastic Markov processes [63]; their ground states are equal-weight superpositions of all the configurations in a sector, and their spectral gaps can be bounded by the Markov-chain gap. The tools developed here may be useful for undertaking a more detailed study of properties of these states. The family of models we considered is restricted in the sense that the dynamical constraints can be expressed in the computational basis, so that every computational-basis product state is in a definite dynamical sector of Hilbert space. More generally, one can consider constraints associated with a commuting set of projectors with entangled eigenstates. In one-dimensional spin chains, such commuting projectors can be deformed into unentangled projectors by a short-depth unitary circuit. However, this process just corresponds to conjugating the Hamiltonians/unitaries we have explored with short-depth unitary circuits, and yields nothing qualitatively new. Getting something new in one dimension thus requires defining constraints using non-commuting projectors—which occurs in the (quantum-fragmented) Temperley-Lieb model [64, 65, 23]—and hence pursuing this direction requires systematic characterization of such constraints. In higher dimensions however, there are sets of commuting projectors (like those of the toric code) whose simultaneous eigenstates are inherently long-range entangled. An interesting task for future work would be to extend our group-theoretic dynamical systems to these models, and explore the resulting entanglement dynamics. Finally, it would be interesting to explore the stability of our results with respect to weak violations of the constraints. One way to do this is by breaking the constraints in an isolated region of space, or by bringing a thermal bath in contact with the system at its boundary. In the case of group dynamics, doing so destroys fragmentation, making the dynamics fully ergodic. It furthermore leads to all states being connected by at most $O(L^{2})$ steps of the resulting dynamics, since the $O(L)$ elements of a given product state’s geodesic word need to be moved at most an $O(L)$ distance to the constraint-free region, at which point they can be changed into the elements of any other geodesic word. It is however still possible for the dynamics in this case to have long (even exponentially long in $L$) thermalization times, due to bottlenecks in Fock space that arise from the finite spatial extent of the constraint-free region. A specific example of this is presented in [66].
@@ IX Acknowledgements
Appendix A Semigroup presentations
In this section we review some basic notions about discrete semigroup presentations and prove a few small results mentioned in the main text. Formally, a discrete semigroup $G$ is a set equipped with a binary associative operation; when $G$ is a group it additionally has a distinguished element that acts as the identity, and each member of the set has a corresponding inverse. It is common to discuss a semigroup $G$ in terms of a specific set of generators $S$ and relations $R$ between them, writing
$G={\sf semi}\langle S\,|\,R\rangle$ (119)
to signify this relationship (in the main text, when $G$ is a group we will write simply $G=\langle S\,|\,R\rangle$ and omit inverse generators and the identity from $S$, as well as trivial relations involving the identity from $R$). For example, one might think of the group $\mathbb{Z}$ as being defined by a single generator $S=\{\mathtt{x}\}$ which obeys no nontrivial relations. This is however too narrow of a viewpoint, since it is possible for different choices of $S$ and $R$ to produce the same semigroup. As an example, consider the groups
@@ Appendix A Semigroup presentations
$G_{2}$ $=\langle\mathtt{x},\mathtt{y}\,|\,\mathtt{x}^{m}=\mathtt{y}^{n},\,\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}\rangle$
where $m,n$ are relatively prime. These two groups are isomorphic, $G_{1}\cong G_{2}\cong\mathbb{Z}$, with the isomorphism associating an element $\mathtt{x}^{a}\mathtt{y}^{b}$ with the integer $am+bn$; note that this is true despite the fact that $G_{1}$ and $G_{2}$ have a different number of generators and relations. A given semigroup in general admits an infinite number of different presentations, but below we will prove that the group-theoretic functions defined in the main text—the Dehn function, expansion length, and so on—have asymptotic scaling behaviors that are presentation independent. We can exploit this fact to choose presentations satisfying some particular desired property. For example, we may be concerned with choosing a model of dynamics where the Hamiltonian or unitary gates are as local as possible. Since the locality of the dynamics is limited by the maximum size of the relations in $R$, we would thus like to minimize the size of the relations. To this end we have the following proposition:
@@
Every finitely generated group has a presentation $\langle S|R\rangle$ where all relations $r_{i}\in R$ have length $|r_{i}|\leq 3$.
@@
Consider the Cayley 2-complex ${\mathsf{CG}}_{G}$ of a finitely presentable group $G$ (see Sec. IV for a brief discussion of its definition) obtained from a finite presentation $P$. While the exact structure of ${\mathsf{CG}}_{G}$ depends on $P$, ${\mathsf{CG}}_{G}$ can always be subdivided to obtain a simplicial complex where each 2-cell has 3 edges. Since each 2-cell in the complex corresponds to a relation, $|r_{i}|\leq 3$ for all $r_{i}\in R$ in the subdivided complex, thereby defining a presentation $P^{\prime}$ whose relations all have length $\leq 3$. Since $P$ is finite this subdivision is completed after only a finite number of steps, and $P^{\prime}$ is consequently also finite. ∎
Note that while the Cayley 2-complex of a semigroup can also be subdivided, the lack of translation invariance in a semigroup’s Cayley complex means that the resulting subdivision may yield a presentation with infinitely many generators (an illustrative example is to compare the semigroup $\mathbb{N}\times\mathbb{N}$ with the group $\mathbb{Z}\times\mathbb{Z}$). As mentioned above, the group theoretic properties we are interested in from the point of view of group dynamics are largely insensitive to the choice of presentation. For example, recall the growth function $N_{K}(L)\triangleq|\{g,\,|g|\leq L\}|$ defined in (30) of the main text, which measures the number of dynamical sectors as a function of system size. The scaling of $N_{K}(L)$ with $L$ is independent of the choice of presentation, allowing us to meaningfully talk about the growth function of a semigroup, rather than of a presentation:
@@
Let $N_{K,P}$ denote the growth function for a particular presentation $P={\sf semi}\langle S|R\rangle$ of $G$. Then
@@
Let $P={\sf semi}\langle a_{1},\dots,a_{|S|}\,|\,R\rangle$ and $P^{\prime}={\sf semi}\langle a_{1}^{\prime},\dots,a^{\prime}_{|S^{\prime}|}\,|\,R^{\prime}\rangle$. Then since both $P,P^{\prime}$ present $G$, each $a_{i}^{\prime}$ can be expressed as a product of a finite number of $a_{i}$. Let $n_{PP^{\prime}}$ denote the maximal number of generators of $P$ that appear when writing the $a_{i}^{\prime}$ in terms of these generators. Let also $|g|_{P}$ denote the geodesic distance of $g\in G$ with respect to the presentation $P$. Then $|g|_{P^{\prime}}\leq n_{PP^{\prime}}|g|_{P}$. Thus
@@
We may also perform a similar rewriting of generators of $P$ in terms of those of $P^{\prime}$. After running the same argument, we find that there exist $O(1)$ constants $n_{PP^{\prime}},n_{P^{\prime}P}$ such that
@@
Similar reasoning can be applied to show that the Dehn function and expansion length (see Sec. II or the following appendix for definitions) of a semigroup are presentation-independent:
@@
Let ${\sf Dehn}_{P}(L)$, ${\sf EL}_{P}(L)$ be the Dehn function and expansion length of a semigroup with a finite presentation $P=\langle S|R\rangle$. Then if $P,P^{\prime}$ are any two such finite presentations,
@@ Appendix B A primer in geometric group theory
In this appendix we state and prove some useful facts about the geometry and complexity of finitely presentable discrete groups. Many of the statements derived below are well-known results in the math literature, and we have tried to provide citations to the original works when appropriate. A particularly accessible review of background material relevant to the discussion to follow can be found in Ref. [41]; a more advanced reference is Ref. [67]. As a small notational convenience, in the following the notation $w\sim g$ will be used as shorthand to denote that the word $w$ evalutes to $g$; in the main text this was written as $\varphi(|w\rangle)=g$:
$w\sim g\quad\leftrightarrow\quad\varphi(|w\rangle)=g.$ (125)
We will mostly be interested in infinite groups, since finite ones have trivial large-scale geometry (in a sense soon to be made precise). Finitely presentable infinite semigroups are of course very easily constructed; indeed it is easily verified any group presentation where the number of generators exceeds the number of nontrivial relations [^26] By a “trivial relation” we mean a free reduction/expansion, namely a relation of the form $\mathtt{a}\mathtt{a}^{-1}=\mathtt{e}$. will generate an infinite group. Free groups (on $n>1$ generators) and Abelian groups in some sense define opposite extremes, since the free group has a Cayley graph that is embeddable in the hyperbolic space $\mathbb{H}^{n}$, while Abelian groups have Cayley graphs that are embeddable in $\mathbb{R}^{n}$. Most of the interesting cases for us correspond to when an intermediate amount of ‘Abelian-ness’ is introduced to the non-Abelian free group.
@@
The definition of the Dehn function (14) in the main text relates only to the (worst-case) complexity of deforming a given word $w\in K_{e}$ into the identity word (recall that $K_{g,L}$ is the set of length-$L$ words that represent the element $g$, namely $K_{g}=\{w\,|\,|w|=L,g(w)=g\}$). We claimed in the main text that focusing on the complexity of words in $K_{e}$—as opposed to $K_{g}$ for $g\neq e$—was done without loss of generality. We now prove that studying the complexity of the word problem in $K_{g\neq e}$ does indeed not produce anything which is not already captured by ${\sf Dehn}(L)$:
@@
where $\lesssim$ denotes equivalence up additional contributions linear in $L$ [^27] These additional contributions occur because we need $O(L)$ steps just to get rid of trivial words like $ww^{-1}$. However, an extra $L$ steps can be important when we do not allow words of length greater than $L$ to appear at any stage of the deformation (which we very well may want to do on physical grounds). This means that ${\sf Dehn}(L)\lesssim{\sf Dehn}_{g}(L)$. Conversely, since we can deform $e$ to $w_{1}^{-1}w_{2}\sim e$ in time $\sim{\sf Dehn}(w_{1}w_{2}^{-1})$, $w_{1}$ can be deformed into $w_{1}(w_{1}^{-1}w_{2})=w_{2}$ in time $\lesssim{\sf Dehn}(L)$. Thus we also have ${\sf Dehn}(w,w^{\prime})\lesssim{\sf Dehn}(w_{1}w_{2}^{-1})$, so up to factors of order $L$, we have
@@
The above argument means that—if the amount of space available to us is not restricted—all sectors have asymptotically the same worst-case complexity. If we however require that all words involved have length $\leq L$, Prop. 4 is modified to read
@@
In this case, all sectors $K_{g,L}$ for which the geodesic length [^28] Recall that the geodesic length of a group element is defined by the length of the shortest word representing $g$, namely $|g|=\min\{w\,:\,\varphi(|w\rangle)=g\}$. of $g$ satisfies $|g|/L<1$ in the $L\rightarrow\infty$ limit will still have ${\sf Dehn}_{g}(L)\sim{\sf Dehn}(L)$. On the other hand, sectors where $|g|/L=1$ as $L\rightarrow\infty$ will not have enough “free space” for the above argument to work, and will thus have ${\sf Dehn}_{g}(L)<{\sf Dehn}(L)$—each word in this sector must have at least $|g|$ locations allotted to represent $g$, giving it less room to form large loops in the Cayley graph with the remaining locations (in the case where fragile fragmentation occurs [see Sec. V], this remains true provided we define ${\sf Dehn}(w,w^{\prime})=0$ if $w_{1},w_{2}$ are in different subsectors of $K_{g,L}$). Sectors where $|g|/L=1$ as $L\rightarrow\infty$ are however necessarily exponentially smaller in size than those with $|g|/L<1$ (see Prop. 10), and hence a random state—which with high probability is associated to a geodesic word where $L-|g|$ diverges as $L\rightarrow\infty$—will be in a sector with ${\sf Dehn}_{g}(L)\sim{\sf Dehn}(L)$ with high probability. We now provide examples of simple groups and their Dehn functions as well as discuss which sorts of groups we can expect to have large Dehn functions. A basic result is that only infinite non-Abelian groups can have interesting Dehn functions, due to the following easily verified statements:
@@
Free groups have ${\sf Dehn}(L)\sim L^{2}$.
@@
This follows from the fact that the Cayley graphs of free groups are trees, and thus any word $w\in K_{e}$ is a (potentially backtracking) path on the tree which contains no nontrivial loops. Such paths can be contracted to the trivial path by applying $O(L)$ relations $\mathtt{x}\mathtt{x}^{-1}=\mathtt{e}\mathtt{e}$, and $O(L^{2})$ relations $\mathtt{x}\mathtt{e}=\mathtt{e}\mathtt{x}$ [^31] In the mathematical literature, a convention is often used where ${\sf Dehn}(w)$ is the number of relations needed to reduce $w$ to the identity $\mathtt{e}$, allowing the length of the word to change in the process (so that e.g. $\mathtt{e}\mathtt{e}$ may be replaced by $\mathtt{e}$). For physical models of group dynamics, where the total length of the word is always fixed, ${\sf Dehn}(w)$ will generically be longer than the value obtained with this convention by a factor of $L$, corresponding to moves of the form $\mathtt{e}\mathtt{a}=\mathtt{a}\mathtt{e}$ that need to be performed during the reduction of $w$ to $\mathtt{e}^{L}$.
To look for interesting Dehn functions we thus need infinite groups with a moderate amount of Abelianness, which (roughly speaking) possess nontrivial loops at all length scales. One can start by finding examples of groups where ${\sf Dehn}(L)$ scales as a higher order polynomial. It was shown in Ref. [42] that these turn out to be virtually nilpotent groups. The simplest example is:
@@
has ${\sf Dehn}(L)\sim L^{3}$ [42].
The group ${\sf BS}(1,2)$ [68] studied in detail in the main text is the simplest group known to the authors with exponential Dehn function, whose properties we discuss in detail in App. E.
@@
As discussed in the main text, the space complexity of the word problem is given by the maximal size of words that one must encounter when reducing a $w\in K_{e}$ to the identity, as measured by the expansion length ${\sf EL}(w)$, a quantity originally introduced by Gromov in Ref. [69]. If ${\sf EL}(w)>|w|$, then $w$ must expand by a nontrivial amount while being reduced to the identity. As we did with the area, we may use the expansion length to define a distance metric between any two words $w_{1,2}$ that represent the same group element. Instead of using ${\sf EL}(w_{1}w_{2}^{-1})$ (which we do not do on account of the homotopy relating $w_{1}$ to $w_{2}$ needing to be properly based), we define the relative expansion length ${\sf EL}(w,w^{\prime})$ between two words $w_{1}\sim w_{2}$ by replacing $e$ by $w_{2}$ in the above definition: